Suppose E and F are events in a sample space. Given that Pr(E)=2/3 and Pr(F)=1/3, if Pr (E|F)= 1/2, calculate Pr(F|E).
I tried to solve this problem and got the answer 1/4 for Pr (F|E). Let me know if this is right!! Thank you

Answer by stanbon(75887)

Suppose E and F are events in a sample space. Given that Pr(E)=2/3 and Pr(F)=1/3, if Pr (E|F)= 1/2, calculate Pr(F|E).
I tried to solve this problem and got the answer 1/4 for Pr (F|E). Let me know if this is right!!
-----
Pr(E)=2/3 and Pr(F)=1/3, if Pr (E|F)= 1/2
P(E|F) = P(F and E)/P(F) = 1/2
So, P(F and E) = (1/2)*P(F) = (1/2)(1/3) = 1/6
--------
Your Problem::
P(F|E) = P(F and E)/P(E) = (1/6)/(2/3) = 1/4
-----------------------------------
You are correct.
----
Cheers,
Stan H.
------